OpenStudy (anonymous):

I'm literally about to cry from this problem, medal and fan will be rewarded for help. :( Linda is studying the sale of a particular brand of cereals from the year 1993 to 2004. She writes the following function to model the sale of the cereal S(t), in million dollars, after t years: S(t) = t^2 + 5t + 52 Part A: What does the y-intercept of the graph of the function represent? Part B: What is the reasonable domain of the graph of the function? Part C: What is the average rate of change of the sale of the cereal from the first year to the fourth year? Show your work.

OpenStudy (anonymous):

do you mean t^2?

OpenStudy (anonymous):

Yeah, I should edit that. :)

OpenStudy (anonymous):

\[S(t) = t^2 + 5t + 52\]right?

OpenStudy (anonymous):

Correct @satellite73

OpenStudy (anonymous):

Is Part A y is zero?

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

Ok, I think it's 52, for sure.

OpenStudy (anonymous):

part A is asking for the domain of the function

OpenStudy (anonymous):

Oh!

OpenStudy (anonymous):

oops that is part B

OpenStudy (anonymous):

Haha, it's cool. XD I think y-intercept is 52.

OpenStudy (anonymous):

A is y intercept, which is what you get if \(t=0\) namely \(52\)

OpenStudy (anonymous):

So, Part A is 52? :)

OpenStudy (anonymous):

i thought part A would be the year 1993 but I'm not sure :/

OpenStudy (anonymous):

yes, and it means that in 1993 they solve $52 million

OpenStudy (anonymous):

Oh! Ok, would you mind helping me with Part B and C? I understand Part A. now :-)

OpenStudy (anonymous):

sure the question is a little fuzzy i suppose it to mean that it is \(S(t) = t^2 + 5t + 52\) where we start counting time in \(1993\)

OpenStudy (anonymous):

Is this for Part B.?

OpenStudy (anonymous):

and stop in \(2004\) since \(2004-1993=11\) i would say the domain is \(0\leq t\leq 11\)

OpenStudy (anonymous):

that would be part B yes

OpenStudy (anonymous):

unless i am reading it wrong and you are supposed to use \(1993\leq t\leq 2004\) but that doesn't make any sense for the first problem, so lets go with the first answer

OpenStudy (anonymous):

Domain for Part B. is \[0 \le t \le 11\]

OpenStudy (anonymous):

Ok, may I ask why the 11 is here? :)

OpenStudy (anonymous):

because it is eleven years from 1993 to 2004

OpenStudy (anonymous):

OOOOoohhhhh duuuurrrr!!! Brain fart, haha, ok. XD

OpenStudy (anonymous):

I understand, so the first one with the eleven is more reasonable as an answer than the 1993 is less than or equal to, etc. etc. ?

OpenStudy (anonymous):

i think so, yes the question is not written very well

OpenStudy (anonymous):

I know, I completely understand.

OpenStudy (anonymous):

What about part C.? Would you mind explaining to me about that one?

OpenStudy (anonymous):

Good morrow sir @halorazer

OpenStudy (anonymous):

Oh! Hold on!

OpenStudy (anonymous):

again it is not written very well is the first year \(t=1\) or is it \(1993\) which would make \(t=0\)?

OpenStudy (anonymous):

\[\frac{ y^2 - y^1 }{ x^2 - x^1 }\]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

i would say \(t=0\) and \(t=3\) i.e. 1993 and 1997

OpenStudy (anonymous):

yes right

OpenStudy (anonymous):

It would look like this:

OpenStudy (anonymous):

i meant \(t=4\)but you have the right idea

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

Wait, hold on.

OpenStudy (anonymous):

\[\frac{s(4)-s(0)}{4}\]

OpenStudy (anonymous):

\[\frac{ 1997 - 1993 }{ 4 - 1 } = \frac{ 4 }{ 3 } = 1.33333...\]

OpenStudy (anonymous):

@Indigo

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@satellite73 help!

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@satellite73 please don't leave!!!

OpenStudy (anonymous):

@satellite73 I still need help!